# Quantitative Propagation of Chaos for 2D Viscous Vortex Model with General Circulations on the Whole Space

### v1

@@ 2.4. Local Relative Entropy Control
In this subsection we prove our second main estimate in the right-handside of Lemma 2.2, say the time derivative of the local relative entropy. Thanks again to the non-positive dissipation term, we only need to control the rest terms, which we recall
@@
$\displaystyle-\frac{1}{k}\sum_{i=1}^{k}\frac{N-k}{N}\int_{({\mathbb{R}}\times{\mathbb{R}}^{2})^{k+1}}G^{N,k+1}\Big{(}m_{k+1}K(x_{i}-x_{k+1})-K\ast\omega(x_{i})\Big{)}\cdot\nabla\log g(m_{i},x_{i})\,\mathrm{d}Z^{k+1}$
$\displaystyle+\frac{1}{k}\sum_{i=1}^{k}\frac{N-k}{N}\int_{({\mathbb{R}}\times{\mathbb{R}}^{2})^{k+1}}G^{N,k+1}\Big{(}m_{k+1}K(x_{i}-x_{k+1})\cdot\nabla_{x_{i}}\log G^{N,k}\Big{)}\,\mathrm{d}Z^{k+1}.$
Step 1: Estimate of $J_{1}$. The first self-contained term $J_{1}$ is similar as the quantity that we estimate in the previous subsection. Using again the integration by parts with $K=\nabla\cdot V$ where
@@
$\displaystyle\frac{1}{kN}\sum_{i,j=1}^{k}\int_{({\mathbb{R}}\times{\mathbb{R}}^{2})^{k}}\Big{(}m_{j}V(x_{i}-x_{j})-V\ast\omega(x_{i})\Big{)}:\Big{(}\nabla_{x_{i}}g^{k}\otimes\nabla_{x_{i}}\frac{G^{N,k}}{g^{k}}\Big{)}\,\mathrm{d}Z^{k}$
$\displaystyle$++ $\displaystyle\frac{1}{kN}\sum_{i,j=1}^{k}\int_{({\mathbb{R}}\times{\mathbb{R}}^{2})^{k}}G^{N,k}\Big{(}m_{j}V(x_{i}-x_{j})-V\ast\omega(x_{i})\Big{)}:\frac{\nabla_{x_{i}}^{2}g^{k}}{g^{k}}\,\mathrm{d}Z^{k}.$
@@
One is easy to observe that, once we extract the factor $k/N$ from the above expression, the remaining quantity has exactly the same structure with the time derivative of the global relative entropy appearing in the previous subsection, see (2.7) and (2.8). Hence we shall simply copy the same estimate, but only make a small scaling of $k/N$ on every $\eta(t)$ appearing before, to control $J_{1}$ by
$\displaystyle\frac{\sigma}{4}I_{k}(G^{N,k}|g^{k})+\frac{k}{N}\cdot C\frac{1+\log(1+t)}{1+t}\frac{N}{k}\Big{(}H_{k}(G^{N,k}|g^{k})+\frac{1}{k}\frac{k^{2}}{N^{2}}\Big{)}$
$\displaystyle$== $\displaystyle\,\frac{\sigma}{4}I_{k}(G^{N,k}|g^{k})+C\frac{1+\log(1+t)}{1+t}\Big{(}H_{k}(G^{N,k}|g^{k})+\frac{k}{N^{2}}\Big{)},$
@@
where in the second term of the first line, the factor $N/k$ is due to the scaling of $\eta(t)$ and the factor $k^{2}/N^{2}$ is due to extended Large Deviation Theorem 2.4, where the upper bound $\gamma$ is rescaled corresponding to $\eta(t)$.
@@
Step 2: Estimate of $J_{2}+J_{3}$. As for the rest term $J_{2}+J_{3}$, which is no longer self-contained since the integral includes the $(k+1)$-marginal function, we adapt the idea of [Wan24]. First we use the symmetry property and regroup the terms into
$\displaystyle\frac{N-k}{N}\int_{({\mathbb{R}}\times{\mathbb{R}}^{2})^{k+1}}G^{N,k}\Big{(}m_{k+1}K(x_{1}-x_{k+1})\frac{G^{N,k+1}}{G^{N,k}}-K\ast\omega(x_{1})\Big{)}\cdot\nabla_{x_{1}}\log\frac{G^{N,k}}{g^{k}}\,\mathrm{d}Z^{k+1},$
@@
where we also make use of the divergence-free property of the kernel $K$. Now we first perform a Cauchy-Schwartz argument to control by
$\displaystyle\frac{\sigma}{4}\int_{({\mathbb{R}}\times{\mathbb{R}}^{2})^{k}}G^{N,k}\Big{|}\nabla_{x_{1}}\log\frac{G^{N,k}}{g^{k}}\Big{|}^{2}\,\mathrm{d}Z^{k}$
$\displaystyle$++ $\displaystyle\frac{1}{\sigma}\int_{({\mathbb{R}}\times{\mathbb{R}}^{2})^{k+1}}G^{N,k}\Big{|}m_{k+1}K(x_{1}-x_{k+1})\frac{G^{N,k+1}}{G^{N,k}}-K\ast\omega(x_{1})\Big{|}^{2}\,\mathrm{d}Z^{k+1}.$
@@
The first part is nothing but the relative Fisher information and is cancelled again by the non-positive dissipation term in the right-handside of (2.2), hence we only need to focus on the second part, which we shall rewrite in
$\frac{1}{\sigma}\int_{({\mathbb{R}}\times{\mathbb{R}}^{2})^{k}}G^{N,k}\Big{|}\int_{{\mathbb{R}}\times{\mathbb{R}}^{2}}m_{k+1}K(x_{1}-x_{k+1})\Big{(}\frac{G^{N,k+1}}{G^{N,k}}-g(m_{k+1},x_{k+1})\Big{)}\,\mathrm{d}z_{k+1}\Big{|}^{2}\,\mathrm{d}Z^{k}.$ (2.11)
@@
The inside integral is in the form of the difference between two probability density functions testing on some singular function, hence we cannot directly bound it by some total variation norm using the Csiszár-Kullback-Pinsker inequality. However, one can perform integration by parts to eliminate the singularity and make use of the relative Fisher information. This has been observed by [Wan24] requiring $\nabla\log g\in L^{\infty}$, hence makes sense in the torus case. Now we generalize this idea to our setting where the logarithmic gradient growth estimate holds as in Proposition 4.1 below, by introducing the weighted CKP inequality initiated by Bolley-Villani [BV05]. We are coming up with the following transport-type estimate.
@@
$|\langle K,m_{1}-m_{2}\rangle|\leq\|V\|_{\infty}\sqrt{I(m_{1}|m_{2})}+\lambda^{-1}\sqrt{1+\log\int e^{\lambda^{2}\|V\|_{\infty}^{2}|\nabla\log m_{2}|^{2}}\,\mathrm{d}m_{2}}\sqrt{2H(m_{1}|m_{2})}.$
@@
This result is a simple combination of the integration by parts process of [Wan24, Proposition 7] and the weighted CKP inequality, see for example in [BV05] or [Lac23, Section 6].
@@
$\displaystyle\,\frac{1}{\sigma}\int_{({\mathbb{R}}\times{\mathbb{R}}^{2})^{k}}G^{N,k}\Big{|}\int_{{\mathbb{R}}\times{\mathbb{R}}^{2}}m_{k+1}K(x_{1}-x_{k+1})\Big{(}\frac{G^{N,k+1}}{G^{N,k}}-g(m_{k+1},x_{k+1})\Big{)}\,\mathrm{d}z_{k+1}\Big{|}^{2}\,\mathrm{d}Z^{k}$
@@
$\displaystyle\,+2\lambda^{-2}\Big{(}1+\log\int_{{\mathbb{R}}\times{\mathbb{R}}^{2}}e^{\lambda^{2}\|V\|_{\infty}^{2}|\nabla\log g|^{2}}g(z_{k+1})\,\mathrm{d}z_{k+1}\Big{)}\cdot H\Big{(}\frac{G^{N,k+1}}{G^{N,k}}\Big{|}g\Big{)}\Big{]}\,\mathrm{d}Z^{k}.$
One shall notice the chain rule or the towering property of the relative entropy and the relative Fisher information gives that
(×$\displaystyle\int_{({\mathbb{R}}\times{\mathbb{R}}^{2})^{k}}G^{N,k}I\Big{(}\frac{G^{N,k+1}}{G^{N,k}}\Big{|}g\Big{)}\,\mathrm{d}Z^{k}$ $\displaystyle=I_{k+1}(G^{N,k+1}|g^{k+1}),$
(×$\displaystyle\int_{({\mathbb{R}}\times{\mathbb{R}}^{2})^{k}}G^{N,k}H\Big{(}\frac{G^{N,k+1}}{G^{N,k}}\Big{|}g\Big{)}\,\mathrm{d}Z^{k}$ $\displaystyle=(k+1)H_{k+1}(G^{N,k+1}|g^{k+1})-kH_{k}(G^{N,k}|g^{k}),$
where we emphasize that the entropy and Fisher information are normalized quantities with the index. As we have the logarithmic gradient estimate Proposition 4.1 and the Gaussian upper bound Lemma 3.2, we may choose $\lambda=\sqrt{1+t}/C$ to ensure the convergence of the integral inside the logarithm, where $C$ depends only linearly on $\|V\|_{\infty}$ and the initial bounds. Hence we conclude by
$\displaystyle J_{2}+J_{3}$ $\displaystyle\,\leq\frac{\sigma}{4}I_{k}(G^{N,k}|g^{k})+\frac{A^{2}\|V\|_{\infty}^{2}}{\sigma}I_{k+1}(G^{N,k+1}|g^{k+1})$
$\displaystyle\,+C\frac{1+\log(1+t)}{1+t}\Big{(}(k+1)H_{k+1}(G^{N,k+1}|g^{k+1})-kH_{k}(G^{N,k}|g^{k})\Big{)}.$
@@
$\frac{\,\mathrm{d}}{\,\mathrm{d}t}H_{k}\leq$ $\,-\frac{\sigma}{2}I_{k}+\frac{A^{2}\|V\|_{\infty}^{2}}{\sigma}I_{k+1}1_{k<N}+C\frac{1+\log(1+t)}{1+t}H_{k}$ (2.12)
$\,+C\frac{1+\log(1+t)}{1+t}((k+1)H_{k+1}-kH_{k})1_{k<N}+C\frac{k}{N^{2}}\frac{1+\log(1+t)}{1+t}.$
Inspired by [Wan24], we introduce the following new ODE hierarchy proposition to deal with the evolution of the local relative entropy, as an extension to more general growth-in-time functions beyond the previous work [Wan24].
@@
Let $T>0$ and let $x_{k},y_{k}\in\mathcal{C}^{1}([0,T];{\mathbb{R}}_{\geq 0})$ for any $1\leq k\leq N$. Suppose that $x_{k+1}\geq x_{k}$ for all $1\displaystyle 1\leq k\leq N-$ 1. Assume that there exists real numbers $c_{1}>c_{2}\geq 0$ and some non-negative time decay function $h(t)$ such that for all $t\in[0,T]$,
@@
The proof of Proposition 2.1 is rather technical and will be presented in the next subsection. We keep on finishing the main proof admitting the proposition above. Under the assumption for the viscosity constant $\sigma$ that $\sigma>\sqrt{2}A\|V\|_{\infty}$, we can apply such ODE hierarchy result for
$x_{k}(t)=kH_{k}(t),\quad y_{k}(t)=kI_{k}(t),\quad c_{1}=\frac{\sigma}{2},\quad c_{2}=\frac{A^{2}\|V\|_{\infty}^{2}}{\sigma},\quad h(t)=C\frac{1+\log(1+t)}{1+t},$
to arrive our desired main result, say
$H_{k}(t)\leq Me^{M\log^{2}(1+t)}\frac{k}{N^{2}}.$
@@ 2.5. Solving the New ODE Hierarchy
The ODE hierarchy Proposition 2.1 is somehow an extension of the similar result in [Wan24, Proposition 5], where the coefficients appearing in the ODE are either constants or exponential functions in time $t$, hence allowing S. Wang to apply the iterated integral estimates in [Lac23]. However, in our cases, the growth in time function is more complicated from the logarithmic estimates, which forces us to deal with more general iterated integrals. The first half of the proof is similar to [Wan24], but in the rest part we shall make a clearer computation.
Step 1: Change of Variables. Inspired by [Wan24], we take the index
@@
which exists since $\lim_{i\rightarrow\infty}i^{5}/(i+1)^{5}=1>c_{2}/c_{1}$. We induce the following new variable $z_{k}(t)$ as a weighted mixture of $x_{i}(t)$ for all $i\geq k$.
@@
Repeating the estimates of [Wan24], we come up with that for any $k\leq N/2$,
@@
where $f(t)=Ch(t)$ for some universal constant $C$. Also the initial data satisfy that for any $k\leq N/2$,
@@
In order to cancel the first term $f(t)z_{k}$ in the right-hand side, we introduce the growth-in-time function
@@
$w_{k}(t)\leq e^{-k\varphi(t)}w_{k}(0)+\int_{0}^{t}e^{-k\big{(}\varphi(t)-\varphi(s)\big{)}}f(s)\Big{(}\frac{k^{2}}{N^{2}}+kw_{k+1}(s)\Big{)}\,\mathrm{d}s.$
@@
$\displaystyle w_{k}(t)$ $\displaystyle\leq e^{-k\varphi(t)}w_{k}(0)+\frac{k^{2}}{N^{2}}\int_{0}^{t}e^{-k\big{(}\varphi(t)-\varphi(s)\big{)}}f(s)\,\mathrm{d}s$
$\displaystyle+k\int_{0}^{t}e^{-k\big{(}\varphi(t)-\varphi(s)\big{)}}f(s)e^{-(k+1)\varphi(s)}w_{k+1}(0)\,\mathrm{d}s$
$\displaystyle+k\frac{(k+1)^{2}}{N^{2}}\int_{0}^{t}e^{-k\big{(}\varphi(t)-\varphi(s)\big{)}}f(s)\int_{0}^{s}e^{-(k+1)\big{(}\varphi(s)-\varphi(r)\big{)}}f(r)\,\mathrm{d}r\,\mathrm{d}s$
$\displaystyle+k(k+1)\int_{0}^{t}e^{-k\big{(}\varphi(t)-\varphi(s)\big{)}}f(s)\int_{0}^{s}e^{-(k+1)\big{(}\varphi(s)-\varphi(r)\big{)}}f(r)w_{k+2}(r)\,\mathrm{d}r\,\mathrm{d}s.$
By simple iteration of the above formula, we are now coming up with the estimate of $w_{k}(t)$ with the iterated integrals, i.e.
@@
$A_{k}^{l}(t_{k})=\Big{(}k(k+1)...l\Big{)}\int_{0}^{t_{k}}...\int_{0}^{t_{l}}$ $\exp\Big{(}-\sum_{j=k}^{l}j\big{(}\varphi(t_{j})-\varphi(t_{j+1})\big{)}\Big{)}$ (2.13)
$f(t_{k+1})...f(t_{l+1})\,\mathrm{d}t_{l+1}...\,\mathrm{d}t_{k+1},$
$B_{k}^{l}(t_{k})=\Big{(}k(k+1)...(l-1)\Big{)}\int_{0}^{t_{k}}...\int_{0}^{t_{l-1}}$ $\exp\Big{(}-\sum_{j=k}^{l-1}j\big{(}\varphi(t_{j})-\varphi(t_{j+1})\big{)}-l\varphi(t_{l})\Big{)}$ (2.14)
$f(t_{k+1})...f(t_{l})\,\mathrm{d}t_{l}...\,\mathrm{d}t_{k+1},$
together with $A_{k}^{k}(t)=1-e^{-k\varphi(t)}$ and $B_{k}^{k}(t)=e^{-k\varphi(t)}$. Also we have the initial values
@@
Step 3: Estimate of Iterated Integrals. We highlight that these two iterated integrals above extend the similar expressions studied in [Lac23, Wan24], where the growth-in-time function is in the specific structure
@@
Lemma 2.6 (Calculations of Iterated Integrals).
@@
$\displaystyle$== $\displaystyle\,\Big{(}\prod_{j=k}^{l-1}j\Big{)}\int_{0}^{t_{k}}...\int_{0}^{t_{l-1}}\exp\Big{(}-\sum_{j=k}^{l-1}j\big{(}\varphi(t_{j})-\varphi(t_{j+1})\big{)}-l\varphi(t_{l})\Big{)}\Big{(}-kf(t_{k})f(t_{k+1})...f(t_{l})\Big{)}$
$\displaystyle\,+\Big{(}\prod_{j=k}^{l-1}j\Big{)}\int_{0}^{t_{k+1}}...\int_{0}^{t_{l-1}}\exp\Big{(}-\sum_{j=k}^{l-1}j\big{(}\varphi(t_{j})-\varphi(t_{j+1})\big{)}-l\varphi(t_{l})\Big{)}f(t_{k+1})...f(t_{l})\Big{|}_{t_{k+1}=t_{k}}$
@@
$\displaystyle\,+\Big{(}\prod_{j=k}^{l-1}j\Big{)}\int_{0}^{t_{k+1}}...\int_{0}^{t_{l-1}}\exp\Big{(}-\sum_{j=k+1}^{l-1}j\big{(}\varphi(t_{j})-\varphi(t_{j+1})\big{)}-l\varphi(t_{l})\Big{)}f(t_{k+1})...f(t_{l})\Big{|}_{t_{k+1}=t_{k}}$

### v2

@@ 2.4. Local Relative Entropy Control
In this subsection we prove our second main estimate on the right-hand side of Lemma 2.2, say the time derivative of the local relative entropy. Thanks again to the non-positive dissipation term, we only need to control the rest terms, which we recall
@@
$\displaystyle-\frac{1}{k}\sum_{i=1}^{k}\frac{N-k}{N}\int_{\mathbb{D}^{k+1}}G^{N,k+1}\Big{(}m_{k+1}K(x_{i}-x_{k+1})-K\ast\omega(x_{i})\Big{)}\cdot\nabla\log g(m_{i},x_{i})\,\mathrm{d}Z^{k+1}$
$\displaystyle+\frac{1}{k}\sum_{i=1}^{k}\frac{N-k}{N}\int_{\mathbb{D}^{k+1}}G^{N,k+1}\Big{(}m_{k+1}K(x_{i}-x_{k+1})\cdot\nabla_{x_{i}}\log G^{N,k}\Big{)}\,\mathrm{d}Z^{k+1}.$
Step 1: Estimate of $J_{1}$. The first self-contained term $J_{1}$ is similar to the quantity that we estimate in the previous subsection. Using again the integration by parts with $K=\nabla\cdot V$ where
@@
$\displaystyle\frac{1}{kN}\sum_{i,j=1}^{k}\int_{\mathbb{D}^{k}}\Big{(}m_{j}V(x_{i}-x_{j})-V\ast\omega(x_{i})\Big{)}:\Big{(}\nabla_{x_{i}}g^{k}\otimes\nabla_{x_{i}}\frac{G^{N,k}}{g^{k}}\Big{)}\,\mathrm{d}Z^{k}$
$\displaystyle$++ $\displaystyle\frac{1}{kN}\sum_{i,j=1}^{k}\int_{\mathbb{D}^{k}}G^{N,k}\Big{(}m_{j}V(x_{i}-x_{j})-V\ast\omega(x_{i})\Big{)}:\frac{\nabla_{x_{i}}^{2}g^{k}}{g^{k}}\,\mathrm{d}Z^{k}.$
@@
It is easy to observe that once we extract the factor $k/N$ from the above expression, the remaining quantity has exactly the same structure as the time derivative of the global relative entropy appearing in the previous subsection, see (2.7) and (2.8). Hence we shall simply copy the same estimate, but only make a small scaling of $k/N$ on every $\eta(t)$ appearing before, to control $J_{1}$ by
$\displaystyle\frac{\sigma}{4}I_{k}(G^{N,k}|g^{k})+\frac{k}{N}\cdot\frac{C}{1+t}\frac{N}{k}\Big{(}H_{k}(G^{N,k}|g^{k})+\frac{1}{k}\frac{k^{2}}{N^{2}}\Big{)}$
$\displaystyle$== $\displaystyle\,\frac{\sigma}{4}I_{k}(G^{N,k}|g^{k})+\frac{C}{1+t}\Big{(}H_{k}(G^{N,k}|g^{k})+\frac{k}{N^{2}}\Big{)},$
@@
where in the second term of the first line, the factor $N/k$ is due to the scaling of $\eta(t)$ and the factor $k^{2}/N^{2}$ is due to the extended Large Deviation Theorem 2.4, where the upper bound $\gamma$ is rescaled corresponding to $\eta(t)$.
@@
Step 2: Estimate of $J_{2}+J_{3}$. As for the rest term $J_{2}+J_{3}$, which is no longer self-contained since the integral includes the $(k+1)$-marginal function, we adapt the idea of [55]. First, we use the symmetry property and regroup the terms into
$\displaystyle\frac{N-k}{N}\int_{\mathbb{D}^{k+1}}G^{N,k}\Big{(}m_{k+1}K(x_{1}-x_{k+1})\frac{G^{N,k+1}}{G^{N,k}}-K\ast\omega(x_{1})\Big{)}\cdot\nabla_{x_{1}}\log\frac{G^{N,k}}{g^{k}}\,\mathrm{d}Z^{k+1},$
@@
where we have also used the divergence-free property of the kernel $K$. Now we first perform a Cauchy-Schwarz argument to control by
$\displaystyle\frac{\sigma}{4}\int_{\mathbb{D}^{k}}G^{N,k}\Big{|}\nabla_{x_{1}}\log\frac{G^{N,k}}{g^{k}}\Big{|}^{2}\,\mathrm{d}Z^{k}$
$\displaystyle$++ $\displaystyle\frac{1}{\sigma}\int_{\mathbb{D}^{k+1}}G^{N,k}\Big{|}m_{k+1}K(x_{1}-x_{k+1})\frac{G^{N,k+1}}{G^{N,k}}-K\ast\omega(x_{1})\Big{|}^{2}\,\mathrm{d}Z^{k+1}.$
@@
The first part is exactly the relative Fisher information and will be cancelled again by the non-positive dissipation term on the right-hand side of (2.2). Therefore, we only need to focus on the second part, which we shall rewrite in
$\frac{1}{\sigma}\int_{\mathbb{D}^{k}}G^{N,k}\Big{|}\int_{\mathbb{D}}m_{k+1}K(x_{1}-x_{k+1})\Big{(}\frac{G^{N,k+1}}{G^{N,k}}-g(m_{k+1},x_{k+1})\Big{)}\,\mathrm{d}z_{k+1}\Big{|}^{2}\,\mathrm{d}Z^{k}.$ (2.11)
@@
The inner integral is in the form of the difference between two probability density functions testing on some singular function, which we cannot directly bound by some total variation norm using the CKP inequality. However, we can perform integration by parts to eliminate the singularity and make use of the relative Fisher information to deal with the extra derivatives. This has been observed by [55] under the condition that $\nabla\log g\in L^{\infty}$, which makes sense in the torus case. Now we generalize this idea to our setting, where the logarithmic gradient growth estimate holds as in Proposition 4.1 below, by introducing the weighted CKP inequality initiated by Bolley-Villani [1]. We are coming up with the following transport-type estimate.
@@
$\displaystyle|\langle K,m_{1}-m_{2}\rangle|$ $\displaystyle\,\leq\|V\|_{L^{\infty}}\sqrt{I(m_{1}|m_{2})}$
$\displaystyle\,+\lambda^{-1}\sqrt{1+\log\int e^{\lambda^{2}\|V\|_{L^{\infty}}^{2}|\nabla\log m_{2}|^{2}}\,\mathrm{d}m_{2}}\sqrt{2H(m_{1}|m_{2})}.$
@@
This result is a simple combination of the integration by parts process of [55, Proposition 7] and the weighted CKP inequality, see for example in [1] or [34, Section 6].
@@
$\displaystyle\,\frac{1}{\sigma}\int_{\mathbb{D}^{k}}G^{N,k}\Big{|}\int_{\mathbb{D}}m_{k+1}K(x_{1}-x_{k+1})\Big{(}\frac{G^{N,k+1}}{G^{N,k}}-g(m_{k+1},x_{k+1})\Big{)}\,\mathrm{d}z_{k+1}\Big{|}^{2}\,\mathrm{d}Z^{k}$
@@
$\displaystyle\,+2\lambda^{-2}\Big{(}1+\log\int_{\mathbb{D}}e^{\lambda^{2}\|V\|_{L^{\infty}}^{2}|\nabla\log g|^{2}}g(z_{k+1})\,\mathrm{d}z_{k+1}\Big{)}\cdot H\Big{(}\frac{G^{N,k+1}}{G^{N,k}}\Big{|}g\Big{)}\Big{]}\,\mathrm{d}Z^{k}.$
We recall the classical chain rule or the towering property of the relative entropy and the relative Fisher information:
$\displaystyle\int_{\mathbb{D}^{k}}G^{N,k}I\Big{(}\frac{G^{N,k+1}}{G^{N,k}}\Big{|}g\Big{)}\,\mathrm{d}Z^{k}$ $\displaystyle=I_{k+1}(G^{N,k+1}|g^{k+1}),$
$\displaystyle\int_{\mathbb{D}^{k}}G^{N,k}H\Big{(}\frac{G^{N,k+1}}{G^{N,k}}\Big{|}g\Big{)}\,\mathrm{d}Z^{k}$ $\displaystyle=(k+1)H_{k+1}(G^{N,k+1}|g^{k+1})-kH_{k}(G^{N,k}|g^{k}),$
where we emphasize that the entropy and Fisher information are normalized quantities with the index. As we have the logarithmic gradient estimate Proposition 4.1 and the Gaussian upper bound Lemma 3.2, we may choose $\lambda=\sqrt{1+t}/C$ to ensure the convergence of the integral inside the logarithm, where $C$ depends only linearly on $\|V\|_{L^{\infty}}$ and the initial bounds. Hence we conclude by
$\displaystyle J_{2}+J_{3}$ $\displaystyle\,\leq\frac{\sigma}{4}I_{k}(G^{N,k}|g^{k})+\frac{A^{2}\|V\|_{L^{\infty}}^{2}}{\sigma}I_{k+1}(G^{N,k+1}|g^{k+1})$
$\displaystyle\,+\frac{C}{1+t}\Big{(}(k+1)H_{k+1}(G^{N,k+1}|g^{k+1})-kH_{k}(G^{N,k}|g^{k})\Big{)}.$
@@
$\frac{\,\mathrm{d}}{\,\mathrm{d}t}H_{k}\leq$ $\,-\frac{\sigma}{2}I_{k}+\frac{A^{2}\|V\|_{L^{\infty}}^{2}}{\sigma}I_{k+1}1_{k<N}+\frac{C}{1+t}H_{k}$ (2.12)
$\,+\frac{C}{1+t}((k+1)H_{k+1}-kH_{k})1_{k<N}+\frac{C}{1+t}\frac{k}{N^{2}}.$
Inspired by [55], we introduce the following new ODE hierarchy proposition to deal with the evolution of the local relative entropy, as an extension to more general growth-in-time functions beyond the previous work [55].
@@
Let $T>0$ and let $x_{k},y_{k}\in C^{1}([0,T];[0,\infty))$ for any $1\leq k\leq N$. Suppose that $x_{k+1}\geq x_{k}$ for all $1\displaystyle 1\leq k\leq N-$ 1. Assume that there exists real numbers $c_{1}>c_{2}\geq 0$ and some non-negative time decay function $h(t)$ such that for all $t\in[0,T]$,
@@
The proof of Proposition 2.1 is rather technical and will be presented in the next subsection. We keep on finishing the main proof admitting the above proposition. Under the assumption for the viscosity constant $\sigma$ that $\sigma>\sqrt{2}A\|V\|_{L^{\infty}}$, we can apply this ODE hierarchy result for
$x_{k}(t)=kH_{k}(t),\quad y_{k}(t)=kI_{k}(t),\quad c_{1}=\frac{\sigma}{2},\quad c_{2}=\frac{A^{2}\|V\|_{L^{\infty}}^{2}}{\sigma},\quad h(t)=\frac{C}{1+t},$
to arrive at our desired main result, say
$H_{k}(t)\leq M(1+t)^{M}\frac{k}{N^{2}}.$
@@ 2.5. Solving the New ODE Hierarchy
The ODE hierarchy Proposition 2.1 is somehow an extension of the similar result in [55, Proposition 5], where the coefficients appearing in the ODE are either constants or exponential functions in time $t$, hence allowing S. Wang to apply the iterated integral estimates in [34]. However, in our cases, the growth in time function is more complicated from the logarithmic estimates, which forces us to deal with more general iterated integrals. The first half of the proof is similar to [55], but in the rest part we shall make a clearer computation.
Step 1: Change of Variables. Inspired by [55], we take the index
@@
which exists since $\lim_{i\rightarrow\infty}i^{5}/(i+1)^{5}=1>c_{2}/c_{1}$. We induce the following new variable $z_{k}(t)$ as a weighted mixture of $x_{i}(t)$ for all $i\geq k$:
@@
Repeating the estimates of [55], we arrive at that for any $k\leq N/2$,
@@
where $f(t)=Ch(t)$ for some universal constant $C$. Also, the initial data satisfy for any $k\leq N/2$,
@@
In order to cancel the first term $f(t)z_{k}$ on the right-hand side, we introduce the growth-in-time function
@@
$w_{k}(t)\leq e^{-k\varphi(t)}w_{k}(0)+\int_{0}^{t}e^{-k(\varphi(t)-\varphi(s))}f(s)\Big{(}\frac{k^{2}}{N^{2}}+kw_{k+1}(s)\Big{)}\,\mathrm{d}s.$
@@
$\displaystyle w_{k}(t)$ $\displaystyle\leq e^{-k\varphi(t)}w_{k}(0)+\frac{k^{2}}{N^{2}}\int_{0}^{t}e^{-k(\varphi(t)-\varphi(s))}f(s)\,\mathrm{d}s$
$\displaystyle+k\int_{0}^{t}e^{-k(\varphi(t)-\varphi(s))}f(s)e^{-(k+1)\varphi(s)}w_{k+1}(0)\,\mathrm{d}s$
$\displaystyle+k\frac{(k+1)^{2}}{N^{2}}\int_{0}^{t}e^{-k(\varphi(t)-\varphi(s))}f(s)\int_{0}^{s}e^{-(k+1)(\varphi(s)-\varphi(r))}f(r)\,\mathrm{d}r\,\mathrm{d}s$
$\displaystyle+k(k+1)\int_{0}^{t}e^{-k(\varphi(t)-\varphi(s))}f(s)\int_{0}^{s}e^{-(k+1)(\varphi(s)-\varphi(r))}f(r)w_{k+2}(r)\,\mathrm{d}r\,\mathrm{d}s.$
By simple iteration of the above formula, we are now arriving at the estimate of $w_{k}(t)$ with the iterated integrals, i.e.
@@
$A_{k}^{l}(t_{k})=\Big{(}k(k+1)\cdots l\Big{)}\int_{0}^{t_{k}}\cdots\int_{0}^{t_{l}}$ $\exp\Big{(}-\sum_{j=k}^{l}j\big{(}\varphi(t_{j})-\varphi(t_{j+1})\big{)}\Big{)}$ (2.13)
$f(t_{k+1})\cdots f(t_{l+1})\,\mathrm{d}t_{l+1}\cdots\,\mathrm{d}t_{k+1},$
$B_{k}^{l}(t_{k})=\Big{(}k(k+1)\cdots(l-1)\Big{)}\int_{0}^{t_{k}}\cdots\int_{0}^{t_{l-1}}$ $\exp\Big{(}-\sum_{j=k}^{l-1}j\big{(}\varphi(t_{j})-\varphi(t_{j+1})\big{)}-l\varphi(t_{l})\Big{)}$ (2.14)
$f(t_{k+1})\cdots f(t_{l})\,\mathrm{d}t_{l}\cdots\,\mathrm{d}t_{k+1},$
together with $A_{k}^{k}(t)=1-e^{-k\varphi(t)}$ and $B_{k}^{k}(t)=e^{-k\varphi(t)}$. Also, we have the initial values
@@
Step 3: Estimate of Iterated Integrals. We highlight that these two iterated integrals above extend the similar expressions studied in [34, 55], where the growth-in-time function is in the specific structure
@@
Lemma 2.6 (Computations of Iterated Integrals).
@@
$\displaystyle$== $\displaystyle\,\Big{(}\prod_{j=k}^{l-1}j\Big{)}\int_{0}^{t_{k}}\cdots\int_{0}^{t_{l-1}}\exp\Big{(}-\sum_{j=k}^{l-1}j\big{(}\varphi(t_{j})-\varphi(t_{j+1})\big{)}-l\varphi(t_{l})\Big{)}\Big{(}-kf(t_{k})f(t_{k+1})\cdots f(t_{l})\Big{)}$
$\displaystyle\,+\Big{(}\prod_{j=k}^{l-1}j\Big{)}\int_{0}^{t_{k+1}}\cdots\int_{0}^{t_{l-1}}\exp\Big{(}-\sum_{j=k}^{l-1}j\big{(}\varphi(t_{j})-\varphi(t_{j+1})\big{)}-l\varphi(t_{l})\Big{)}f(t_{k+1})\cdots f(t_{l})\Big{|}_{t_{k+1}=t_{k}}$
@@
$\displaystyle\,+\Big{(}\prod_{j=k}^{l-1}j\Big{)}\int_{0}^{t_{k+1}}\cdots\int_{0}^{t_{l-1}}\exp\Big{(}-\sum_{j=k+1}^{l-1}j\big{(}\varphi(t_{j})-\varphi(t_{j+1})\big{)}-l\varphi(t_{l})\Big{)}f(t_{k+1})\cdots f(t_{l})\Big{|}_{t_{k+1}=t_{k}}$